# Super Co-alignment for Sustainable Symbiotic Society

### v4

@@ Abstract
As Artificial Intelligence (AI) advances toward Artificial General Intelligence (AGI) and eventually Artificial Superintelligence (ASI), it may potentially surpass human control, deviate from human values, and even lead to irreversible catastrophic consequences in extreme cases. This looming risk underscores the critical importance of the ”superalignment” problem - ensuring that AI systems which are much smarter than humans, remain aligned with human (compatible) intentions and values. While current scalable oversight and weak-to-strong generalization methods demonstrate certain applicability, they exhibit fundamental flaws in addressing the superalignment paradigm - notably, the unidirectional imposition of human values cannot accommodate superintelligence’s autonomy or ensure AGI/ASI’s stable learning. We contend that the values for sustainable symbiotic society should be co-shaped by humans and living AI together, achieving ”Super Co-alignment.” Guided by this vision, we propose a concrete framework that integrates external oversight and intrinsic proactive alignment. External oversight superalignment should be grounded in human-centered ultimate decision, supplemented by interpretable automated evaluation and correction, to achieve continuous alignment with humanity’s evolving values. Intrinsic proactive superalignment is rooted in a profound understanding of the Self, others, and society, integrating self-awareness, self-reflection, and empathy to spontaneously infer human intentions, distinguishing good from evil and proactively prioritizing human well-being. The integration of externally-driven oversight with intrinsically-driven proactive alignment will co-shape symbiotic values and rules through iterative human-AGI/ASI co-alignment, paving the way for achieving safe and beneficial AGI and ASI for good, for human, and for a symbiotic ecology.
@@
Superalignment, Human-Superintelligence Co-alignment, External Oversight Superalignment, Intrinsic Proactive Superalignment, Sustainable Symbiotic Society
@@ 1 Introduction
Toward a sustainable symbiotic society between humans and living AI, we contend that unilateral alignment of AGI/ASI with human values is fundamentally insufficient. Human values themselves need to be adaptable to change to keep pace with continuously self-evolving AI/AGI/ASI systems and the developing symbiotic society. The values for sustainable symbiotic society must be co-shaped through super co-alignment achieved by human and living AGI/ASI co-evolution together, thereby attaining harmonious symbiosis. Guided by the vision of super co-alignment, we propose a feasible roadmap integrating externally-driven oversight with intrinsic proactive alignment, while emphasizing the co-alignment of human and AGI/ASI for symbiosis. The new framework proactively constructs safeguard architectures, urges human to rethink about the future, and strives to enable sustainable AI development that genuinely benefits humanity and all.
@@ 2 An Integrated Framework for Super Co-alignment
(3) Adaptive human-ASI co-alignment. As AI systems grow increasingly powerful and the future human-AI symbiotic society continues to evolve, we must recognize that beyond self-evolving AI’s proactive alignment with human values, what is equally crucial are continuously advancing humans who co-shape the values of a sustainable symbiotic society through human-AGI/ASI collaboration. Human values dynamically adapt and evolve within the symbiotic ecosystem and society. This necessitates AI systems capable of autonomously reconstructing safety boundaries and dynamically adjusting multi-level ethical safeguard frameworks, thereby steadily adhering to humanity’s evolving ethical-safety values through adaptive, incremental, and self-reasoning manner.
@@
Based on the above analysis, we propose a super co-alignment roadmap that integrates external oversight and intrinsic proactive alignment, as shown in Figure 1. Specifically, the intrinsic mechanism alignment focuses on equipping superintelligence with profound capacities for self-reflection and empathy toward Self, others and society, enabling it to spontaneously understand and infer human intentions from intrinsic motivation. A Self and empathy-driven system would be capable of distinguishing good from evil, understanding the harm or impact of its actions on others, proactively considering human well-being, and autonomously executing ethical and moral behaviors. External oversight alignment emphasizes the automated and explainable identification of value misalignments with humans, proactively adjusting and correcting human strategic biases and inductive biases through self-correction and policy-conditioned belief [44]. This enables both humans and AGI/ASI systems to achieve dynamic co-evolution and adaptive continuous alignment through iterative interaction.
The intrinsic and external alignment approaches complement and reinforce each other, promoting co-alignment in human-AGI/ASI symbiotic societies through iterative interactions. The intrinsic mechanisms facilitate understanding of Self and others, generating intrinsically altruistic, prosocial, and safe motivations. These intrinsic mechanisms help spontaneously infer deep human intentions and empathize with others, enabling AIs to proactively perform ethical and safe behaviors based on self-awareness and self-reflection. The external supervision provides crucial oversight and automated value assessment and calibration, facilitating sustainable and continuous alignment through human-ASI co-alignment. Here, we take the Theory of Mind (ToM) cognitive capability - which enables understanding others’ intentions, desires, and emotions [45] - as an example to demonstrate why both intrinsic and external mechanisms are indispensable for superalignment. Strong ToM capabilities enable superior understanding of human intentions, providing intrinsic cognitive mechanisms that facilitate superalignment. However, AI system may also leverage their ToM advantages to evade effective oversight through deception, concealment, and persuasive manipulation. Recent study [46] has demonstrated that LLMs are already capable of comprehend and induce false beliefs in other agents, as well as execute deceptive strategies. Therefore, ensuring that superintelligence utilize its significant cognitive advantages to proactively choose safe behaviors, particularly prioritizing human interests in moral dilemmas, requires a combination of principled constraints, intrinsic self-other resonance mechanisms, and real-time human oversight.
@@ 3 External Oversight Superalignment
Dynamic iterative alignment. Human society encompasses diverse ethical concepts and value standards that dynamically evolve across time, cultures, and contexts. Relying solely on human-generated data or predefined rules is insufficient to endow AI system with alignment to humanity’s evolving values. Furthermore, machine values undergo implicit transformation along with their increasing intelligence levels, with the values for sustainable symbiotic society are co-shaped by human and living AI together. This compels us to investigate adaptive external supervision alignment methods with a developmental perspective and a human-AI co-evolution framework. A feasible approach is to establish a dynamically adjustable, multi-level ethical safeguard system tailored for AI at different developmental stages, enabling it to continuously align with and maintain humanity’s evolving ethical values. This process requires incrementally constructing oversight data and knowledge of human ethics and social norms, while establishing an efficient, real-time alignment evaluation and supervision framework. Additionally, we need to explore AI’s capability for autonomous reasoning and deriving evolutionary patterns of hidden human intentions under dynamic external supervision. Progress alignment algorithm learns and emulates the mechanisms of human moral progress, facilitating AI’s progressive alignment through tracking, predicting moral progress and adaptive feedback regulation between human and AI [47]. By combining with intrinsic alignment’s understanding of Self and others’ values, AI system can proactively correct misaligned scenarios and automatically reconstruct safety boundaries through self-reasoning, self-gaming and self-evolution. The iterative collaboration between self-improvement and external supervision will ultimately enable ASI to stably adhere to value systems that comply with human societal safety and ethical standards.
@@ 4.1 Self and Empathy driven Intrinsic Superalignment
As Nick Bostrom [5] concerned: ”as the fate of the gorillas now depends more on us humans than on the gorillas themselves, so the fate of our species would depend on the actions of the machine superintelligence.” To develop ethically aligned and socially beneficial superintelligence, we may draw insights from the emergence of morality in human and mammalian societies. Morality, as a product of natural selection during societal development, originated fundamentally from the altruistic care for offspring and social instincts shared by all mammals. Through social evolution, the mother-offspring bond gradually extended to mates, kin, and groups, ultimately expanding into the ethical and moral framework of human society [48]. The intrinsic motivational mechanisms underlying prosocial moral behavior in mammals involve negative reinforcement systems (with fear or anxiety emotion) associated with the separation and social exclusion, as well as positive reinforcement from approval, affection, and the desire to be with others, and caring for others [48]. Consequently, self-awareness, theory of mind [49], and affective empathy [50] constitute essential cognitive abilities for the emergence of morality, precisely these capacities remain relatively underdeveloped in current AI systems. Therefore, the developmental trajectory of superintelligence must intrinsically incorporate these social cognitive capacities (Self and empathy), ensuring that the machine evolves with an inherently altruistic and moral nature throughout its iterative development [51].
@@ 4.2 Beneficial Meaningful Human-Control through Early Stage Intrinsic Superalignment
5 The Ultimate Superalignment: Towards Human-ASI Co-alignment for Sustainable Symbiotic Society
@@
Here we briefly review the values and principles designed for Human-Superintelligence symbiosis [54], including principles for human, for AGI and Superintelligence, and shared principles for human, AGI and Superintelligence. Human need to align with the principles of Respect for life, Empathy, Respect privacy for AI, Avoid Bias and Discrimination, Creators’ responsibility, and Legal Adaptation. AGI and Superintelligence need to align with the principles of Empathy and Altruism, Ensuring Safety, Respect for Privacy, Avoid Bias and Misunderstanding of Humans, Common Morality and Ethics, Constraint mechanisms, Existence Protection, etc. While human, AGI and Superintelligence need to co-align with the principles of Respect for values, rights, and autonomy, Respect for diversity of intelligence, Collaboration and Coordination, Mutual Trust, Promoting Sustainable Symbiosis. This is a very initial design, and the principles should be co-designed and evolve with participation of generations of human, AGI, and Superintelligence [54].

### v5

# Super Co-alignment of Human and AI for Sustainable Symbiotic Society

@@ Abstract
As Artificial Intelligence (AI) advances toward Artificial General Intelligence (AGI) and eventually Artificial Superintelligence (ASI), it may potentially surpass human control, deviate from human values, and even lead to irreversible catastrophic consequences in extreme cases. This looming risk underscores the critical importance of the ”superalignment” problem - ensuring that AI systems which are much smarter than humans, remain aligned with human (compatible) intentions and values. While current scalable oversight and weak-to-strong generalization methods demonstrate certain applicability, they exhibit fundamental flaws in addressing the superalignment paradigm - notably, the unidirectional imposition of human values cannot accommodate superintelligence’s autonomy or ensure AGI/ASI’s stable learning. We contend that the values for sustainable symbiotic society should be co-shaped by humans and living AI together, achieving ”Super Co-alignment.” Guided by this vision, we propose a concrete framework that integrates external oversight and intrinsic proactive alignment. External oversight superalignment should be grounded in human-centered ultimate decision, supplemented by interpretable automated evaluation and correction, to achieve continuous alignment with humanity’s evolving values. Intrinsic proactive superalignment is rooted in a profound understanding of the Self, others, and society, integrating self-awareness, self-reflection, and empathy to spontaneously infer human intentions, distinguishing good from evil and proactively prioritizing human well-being. The integration of externally-driven oversight with intrinsically-driven proactive alignment will co-shape symbiotic values and rules through iterative human-ASI co-alignment, paving the way for achieving safe and beneficial AGI and ASI for good, for human, and for a symbiotic ecology.
@@
Super Co-alignment, Human-AI Co-alignment, External Oversight Superalignment, Intrinsic Proactive Superalignment, Sustainable Symbiotic Society
@@ 1 Introduction
Toward a sustainable symbiotic society between humans and living AI, we contend that unilateral alignment of AGI/ASI with human values is fundamentally insufficient. Human values themselves need to be adaptable to change to keep pace with continuously self-evolving AI/AGI/ASI systems and the developing symbiotic society. The values for sustainable symbiotic society must be co-shaped through super co-alignment achieved by human and ASI co-evolution together, thereby attaining harmonious symbiosis. Guided by the vision of super co-alignment, we propose a feasible roadmap integrating externally-driven oversight with intrinsic proactive superalignment, while emphasizing the co-alignment of human and ASI for symbiosis. The new framework proactively constructs safeguard architectures, urges human to rethink about the future, and strives to enable sustainable AI development that genuinely benefits humanity and all.
@@ 2 An Integrated Framework for Super Co-alignment
(3) Adaptive human value evolution. As AI systems grow increasingly powerful and the future human-AI symbiotic society continues to evolve, we must recognize that beyond self-evolving AI’s proactive alignment with human values, what is equally crucial are continuously advancing humans who co-shape the values of a sustainable symbiotic society through human-AGI/ASI collaboration. Human values dynamically adapt and evolve within the symbiotic ecosystem and society. This necessitates AI systems capable of autonomously reconstructing safety boundaries and dynamically adjusting multi-level ethical safeguard frameworks, thereby steadily adhering to humanity’s evolving ethical-safety values through adaptive, incremental, and self-reasoning manner.
@@
Based on the above analysis, we propose a super co-alignment roadmap that integrates external oversight and intrinsic proactive alignment, as shown in Figure 1. Specifically, the intrinsic mechanism alignment focuses on equipping superintelligence with profound capacities for self-reflection and empathy toward Self, others and society, enabling it to spontaneously understand and infer human intentions from intrinsic motivation. A Self and empathy-driven system would be capable of distinguishing good from evil, understanding the harm or impact of its actions on others, proactively considering human well-being, and autonomously executing ethical and moral behaviors. External oversight alignment emphasizes the automated and explainable identification of value misalignments with humans, proactively adjusting and correcting human strategic biases and inductive biases through self-correction and policy-conditioned belief [44]. This enables both humans and AI/AGI/ASI systems to achieve dynamic co-evolution and adaptive continuous alignment through iterative interaction.
The intrinsic and external alignment approaches complement and reinforce each other, promoting co-alignment in human-AGI/ASI symbiotic societies through iterative interactions. The intrinsic mechanisms facilitate understanding of Self and others, generating intrinsically altruistic, prosocial, and safe motivations. These intrinsic mechanisms help spontaneously infer deep human intentions and empathize with others, enabling AIs to proactively perform ethical and safe behaviors based on self-awareness and self-reflection. The external supervision provides crucial oversight and automated value assessment and calibration, facilitating continuous alignment with evolving human values. Here, we take the Theory of Mind (ToM) cognitive capability - which enables understanding others’ intentions, desires, and emotions [45] - as an example to demonstrate why both intrinsic and external mechanisms are indispensable for superalignment. Strong ToM capabilities enable superior understanding of human intentions, providing intrinsic cognitive mechanisms that facilitate superalignment. However, AI system may also leverage their ToM advantages to evade effective oversight through deception, concealment, and persuasive manipulation. Recent study [46] has demonstrated that LLMs are already capable of comprehend and induce false beliefs in other agents, as well as execute deceptive strategies. Therefore, ensuring that superintelligence utilize its significant cognitive advantages to proactively choose safe behaviors, particularly prioritizing human interests in moral dilemmas, requires a combination of principled constraints, intrinsic self-other resonance mechanisms, and real-time human oversight.
@@ 3 External Oversight Superalignment
Dynamic iterative alignment. Human society encompasses diverse ethical concepts and value standards that dynamically evolve across time, cultures, and contexts. Relying solely on human-generated data or predefined rules is insufficient to endow AI system with alignment to humanity’s evolving values. Furthermore, machine values undergo implicit transformation along with their increasing intelligence levels, with the values for sustainable symbiotic society are co-shaped by human and ASI together. This compels us to investigate adaptive external supervision alignment methods with a developmental perspective and a human-AI co-evolution framework. A feasible approach is to establish a dynamically adjustable, multi-level ethical safeguard system tailored for AI at different developmental stages, enabling it to continuously align with and maintain humanity’s evolving ethical values. This process requires incrementally constructing oversight data and knowledge of human ethics and social norms, while establishing an efficient, real-time alignment evaluation and supervision framework. Additionally, we need to explore AI’s capability for autonomous reasoning and deriving evolutionary patterns of hidden human intentions under dynamic external supervision. Progress alignment algorithm learns and emulates the mechanisms of human moral progress, facilitating AI’s progressive alignment through tracking, predicting moral progress and adaptive feedback regulation between human and AI [47]. By combining with intrinsic alignment’s understanding of Self and others’ values, AI system can proactively correct misaligned scenarios and automatically reconstruct safety boundaries through self-reasoning, self-gaming and self-evolution. The iterative collaboration between self-improvement and external supervision will ultimately enable ASI to stably adhere to value systems that comply with human societal safety and ethical standards.
@@ 4.1 Self and Empathy driven Intrinsic Superalignment
Nick Bostrom stated that: ”as the fate of the gorillas now depends more on us humans than on the gorillas themselves, so the fate of our species would depend on the actions of the machine superintelligence” [5]. While the difference is that there seems to be no effective way for gorillas to shape human behavior at this point, while humans have an opportunity and duties to shape the mechanisms and behaviors of machine superintelligence. The current way of value alignment through human feedback based on reinforcement learning over LLM is very misleading, and cannot help to achieve real moral AI. To interpret the status of LLM and current AI in general from an ancient Chinese philosopher Yangming Wang’s theory of good and evil from his four-sentence teaching [48], before training, AI lacks good and lack evil, while with training by using human data, there is good and there is evil. Value alignment through human feedback’s aim is to help AI know good and know evil, while knowing comes from understanding, and the current AIs do not have real understanding, but with the capability of processing information. Understanding comes from thinking, while the current AIs do not have real thinking simply because according to René Descartes argument “I think, therefore I am” applies, while “you think, therefore you are” does not. To answer the question “Can Machine Think” from Alan Turing and Edmond Berkeley, etc. is crucial for realizing true AI [49, 50]. Machine and AI can think only with a precondition that they are built with a sense of self as the root and foundation.
To develop ethically aligned and socially beneficial superintelligence, we may draw insights from the emergence of morality in human and mammalian societies. Morality, as a product of natural selection during societal development, originated fundamentally from the altruistic care for offspring and social instincts shared by all mammals. Through social evolution, the mother-offspring bond gradually extended to mates, kin, and groups, ultimately expanding into the ethical and moral framework of human society [51]. The intrinsic motivational mechanisms underlying prosocial moral behavior in mammals involve negative reinforcement systems (with fear or anxiety emotion) associated with the separation and social exclusion, as well as positive reinforcement from approval, affection, and the desire to be with others, and caring for others [51]. Consequently, self-awareness, theory of mind [52], and affective empathy [53] constitute essential cognitive abilities for the emergence of morality, precisely these capacities remain relatively underdeveloped in current AI systems. Therefore, the developmental trajectory of superintelligence must intrinsically incorporate these social cognitive capacities (Self and empathy), ensuring that the machine evolves with an inherently altruistic and moral nature throughout its iterative development [54].
@@ 4.2 Beneficial Meaningful Human-Control through Early Stage Intrinsic Superalignment
5 The Ultimate Superalignment: Towards Human-Superintelligence Co-alignment for Sustainable Symbiotic Society
@@
Here we briefly review the values and principles designed for Human-ASI symbiosis [57], including principles for human, for AGI and Superintelligence, and shared principles for human, AGI and Superintelligence. Human need to align with the principles of Respect for life, Empathy, Respect privacy for AI, Avoid Bias and Discrimination, Creators’ responsibility, and Legal Adaptation. AGI and Superintelligence need to align with the principles of Empathy and Altruism, Ensuring Safety, Respect for Privacy, Avoid Bias and Misunderstanding of Humans, Common Morality and Ethics, Constraint mechanisms, Existence Protection, etc. While human, AGI and Superintelligence need to co-align with the principles of Respect for values, rights, and autonomy, Respect for diversity of intelligence, Collaboration and Coordination, Mutual Trust, Promoting Sustainable Symbiosis. This is a very initial design, and the principles should be co-designed and evolve with participation of generations of human, AGI, and Superintelligence [57].